A dog runs along a path near a square garden with side length 10 m . it runs once around always staying exactly 1 m from the edge of the garden . to the nearest meter , how far does the dog run ?
A) 40 m B) 44 m C) 46 m D) 48 m E) 50 m

Answer by mananth(16946)

the side of square garden = 10m
1 meter away from the garden means it goes 1m away on all sides.
the side increases by 2m .
perimeter =4*L =4*12
=48m